# Tree-Wasserstein Distance for High Dimensional Data with a Latent Feature Hierarchy

### v2

@@ 5 Experimental Results
We apply the proposed method to a synthetic example and to word-document and single-cell RNA-sequencing (scRNA-seq) data. We use Word2vec [Mikolov et al., 2013] as word embedding vectors for the word-document data and Gene2vec [Du et al., 2019] as gene embedding vectors in scRNA-seq experiments. The implementation details are in App. D. Additional experiments are in App. E.
@@ 5.1 Synthetic Word-Document Data Example
We compare the proposed TWD with existing TWDs, including (i) Quadtree [Indyk and Thaper, 2003] that constructs a random tree by a recursive division of hypercubes, (ii) Flowtree [Backurs et al., 2020] that computes the optimal flow of Quadtree, (iii) tree-sliced Wasserstein distance (TSWD) [Le et al., 2019] that computes an average of TWDs by random sampling trees, where the numbers of sampling are set to 1, 5, and 10, (iv) UltraTree [Chen et al., 2024] that constructs an ultrametric tree [Johnson, 1967] by minimizing OT regression cost, (v) weighted cluster TWD (WCTWD) [Yamada et al., 2022], (vi) weighted Quadtree TWD (WQTWD) [Yamada et al., 2022], and their sliced variants (vii) SWCTWD and (viii) SWQTWD [Yamada et al., 2022]. We also compare to TWDs where the trees are constructed by (i) MST [Prim, 1957], (ii) Tree Representation (TR) [Sonthalia and Gilbert, 2020] through a divide-and-conquer approach, (iii) gradient-based hierarchical clustering (HC) in hyperbolic space (gHHC) [Monath et al., 2019], (iv) gradient-based Ultrametric Fitting (UltraFit) [Chierchia and Perret, 2019], and (v) HC by hyperbolic Dasgupta’s cost (HHC) [Chami et al., 2020]. We refer to App. B for a detailed description of baselines and highlight the novel aspects of our TWD.
@@ 5.2 Real-World Word-Document Data Analysis
We demonstrate the power of the proposed TWD in real-world document classification tasks. The latent hierarchical structure of words (features) can be viewed as a node branching into related terms [Gollapalli and Caragea, 2014, Yurochkin et al., 2019, Chang and Blei, 2010], allowing for meaningful comparison between documents in OT distance [Kusner et al., 2015], as it measures how words and their associated meanings are transported from one document to another.
We test four word-document benchmarks in Kusner et al. [2015], including BBCSPORT, TWITTER, CLASSIC, and AMAZON. Detailed descriptions of these datasets are reported in App. D. We compare the proposed TWD to the same competing methods as in Sec. 5.1. In addition, we include the Word Mover’s Distance (WMD) [Kusner et al., 2015] that computes the OT distance between documents using Word2Vec [Mikolov et al., 2013]. A dissimilarity classification based on our TWD and the baselines using the $k$NN classifier is applied. We use cross-validation with five trials, where the dataset is randomly divided into $70\%$ training set and $30\%$ testing set, following previous work in word-document data analysis [Kusner et al., 2015, Huang et al., 2016].
Tab. LABEL:tab:classification presents the document classification accuracy. The results of Quadtree, Flowtree, TSWD, and WMD are taken from Huang et al. [2016], Takezawa et al. [2021]. Our method outperforms all the competing baselines on BBCSPORT, TWITTER, and AMAZON datasets, including the pre-trained WMD that is specifically designed for word-document data. On CLASSIC dataset, our proposed TWD yields the second-best classification accuracy and is competitive with WMD.
Table 1: Document and cell classification accuracy (the highest in bold and the second highest underlined). Results marked with are taken from Huang et al. [2016], Takezawa et al. [2021]. The empirical comparisons to TWD methods demonstrate the advantages of our proposed approach, which effectively incorporates the latent feature hierarchy, leading to improved performance.
@@ 5.3 Single-Cell Gene Expression Data Analysis
We further demonstrate the advantages of our TWD in cell classification tasks on scRNA-seq data. The hierarchical structure of genes in scRNA-seq data [Tanay and Regev, 2017] are essential for analyzing complex cellular processes [Ding and Regev, 2021], especially for computing OT distance [Bellazzi et al., 2021, Tran et al., 2021, Bunne et al., 2023]. For instance, in developmental biology, genes associated with stem cell differentiation follow a hierarchical pattern [Kang et al., 2012]. Revealing the latent gene hierarchy is essential for accurately computing the TWD between cells, as it captures the underlying biological relationships and functional similarities.
We test two scRNA-seq datasets in Dumitrascu et al. [2021], including the Zeisel and CBMC. Detailed descriptions of the datasets are reported in App. D. We compare the proposed TWD to the same baselines as in Sec. 5.2. Instead of WMD, we include the Gene Mover’s Distance (GMD) [Bellazzi et al., 2021], a gene-based OT distance using Gene2Vec [Du et al., 2019]. The cell classification tasks are performed in the same way as in Sec. 5.1 and Sec. 5.2. Tab. LABEL:tab:classification shows the mean and the standard deviation of the cell classification accuracy. Our TWD outperforms the competing methods by a large margin. This suggests that the proposed TWD effectively transforms the latent tree-like relationships in gene space, improving the cell classification accuracy compared to random sampling trees and trees constructed by graph-based and HC methods. Notably, our TWD differentiates the cell types better than pre-trained GMD, which uses gene proximity, while ours is geometrically intrinsic.
@@ Acknowledgments
The work of YEL and RT was supported by the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 802735-ERC-DIFFOP.
@@ A.1 Wasserstein Distance
The application of the Wasserstein distance spans various fields, including enhancing generative models in machine learning [Arjovsky et al., 2017, Kolouri et al., 2018], aligning and comparing color distributions in computer graphics [Solomon et al., 2015, Lavenant et al., 2018], and serving as a robust similarity metric in data analysis [Bellazzi et al., 2021, Kusner et al., 2015, Yair et al., 2017].
Let $\mu_{1}$ and $\mu_{2}$ be two probability distributions on a measurable space $\Omega$ with a ground metric $d$, $\Pi(\mu_{1},\mu_{2})$ be the set of couplings $\pi$ on the product space $\Omega\times\Omega$ such that $\pi(B_{1}\times\Omega)=\mu_{1}(B_{1})$ and $\pi(\Omega\times B_{2})=\mu_{2}(B_{2})$ for any sets $B_{1},B_{2}\subset\Omega$. The Wasserstein distance [Villani, 2009] between $\mu_{1}$ and $\mu_{2}$ is defined by
@@
When replacing the $\inf$ in Eq. (4) with the $\arg\min$, the obtained solution $\pi$ is referred to as the Optimal Transport (OT) plan [Monge, 1781, Kantorovich, 1942]. When computing the Wasserstein distance in Eq. (4) between two discrete probabilities with $m$ points, given by
@@
where $\mathbf{C}\in\mathbb{R}^{m\times m}$ is the ground pairwise distance matrix between the coordinates of the discrete probabilities, it can be computed in $O(m^{3}\log m)$ [Peyré et al., 2019], preventing its use of OT for large-scale datasets [Bonneel et al., 2011]. Therefore, existing works have been proposed to reduce this complexity. Among them, a notable work is the Sinkhorn algorithm [Cuturi, 2013, Chizat et al., 2020], which approximates the OT in quadratic time.
@@ A.2 Diffusion Geometry
Diffusion geometry [Coifman and Lafon, 2006] is a mathematical framework designed to analyze high-dimensional data points $\{\mathbf{a}_{j}\in\mathbb{R}^{n}\}_{j=1}^{m}$ by effectively capturing their underlying geometric structures [Belkin and Niyogi, 2008]. It focuses on analyzing the similarity between data points through the process of diffusion propagation. Specifically, the constructed diffusion operator $\mathbf{P}\in\mathbb{R}^{m\times m}$ derived from the observational data is associated with the heat kernel and the Laplacian of the underlying manifold [Belkin and Niyogi, 2008]. Let $\mathbf{Q}(j,j^{\prime})=\exp\left(-d^{2}(j,j^{\prime})/\epsilon\right)$ be the $(j,j^{\prime})$th element of the pairwise affinity matrix between the data points. In the limit $m\rightarrow\infty$ and $\epsilon\rightarrow 0$, the diffusion operator $\mathbf{P}$ converges to the heat kernel $\exp(-\Delta)$, where $\Delta$ is the Laplace–Beltrami operator on the manifold [Coifman and Lafon, 2006]. By applying eigendecomposition to the diffusion operator $\mathbf{P}$, we obtain a sequence of eigenpairs $\{\nu_{j},\bm{\varphi}_{j}\}_{j=1}^{m}$ with positive decreasing eigenvalues $01=\nu_{1}\geq\nu_{2}\geq\ldots\geq\nu_{m}>$ 0. The diffusion maps (DM) in $m^{\prime}<<m$ dimensions at the diffusion time $t$ embeds a point $\mathbf{a}_{j}$ into $\mathbb{R}^{m^{\prime}}$ by
@@
The diffusion operator has been shown to exhibit favorable convergence properties [Coifman and Lafon, 2006]. In the limits, as the number of features $m$ increases toward infinity, the scaling parameter $\epsilon$ approaches zero, and the operator $\mathbf{P}^{t/\epsilon}$ converges pointwise to the Neumann heat kernel $\mathbf{H}_{t}=\exp(-t\Delta)$, associated with the underlying manifold, where $\Delta$ is the Laplace–Beltrami operator on the manifold. That is, the diffusion operator serves as a discrete approximation of the continuous heat kernel, thereby capturing the geometric structure of the manifold within a finite-dimensional framework [Coifman and Lafon, 2006, Singer, 2006, Belkin and Niyogi, 2008].
Note that the density-normalized affinity matrix was considered to mitigate the effect of non-uniform data sampling, which was proposed and shown theoretically in Coifman and Lafon [2006]. In Section 3.1 in Coifman and Lafon [2006], the construction of a family of diffusion operators is introduced with a parameter $\alpha\in\mathbb{R}$ and the density $q$. In Section 3.4 in Coifman and Lafon [2006], it is shown that when $\alpha=1$ (in our case $\widehat{\mathbf{Q}}=\mathbf{D}^{-1}\mathbf{Q}\mathbf{D}^{-1}$), the diffusion operator approximates the heat kernel, and this approximation is independent of the density $q$. This technique has been further explored and utilized in [Mishne et al., 2019, Katz et al., 2020, Shnitzer et al., 2022, 2024].
Diffusion geometry has mainly been utilized in manifold learning, leading to the development of multi-scale, low-dimensional representations and informative distances [Nadler et al., 2005, Coifman and Maggioni, 2006]. Its utility has been demonstrated across various applications spanning various fields, including nonlinear stochastic dynamical systems [Shnitzer et al., 2020], multi-view dimensionality reduction [Lindenbaum et al., 2020b], shape recognition [Bronstein et al., 2010], wavelet analysis [Coifman and Maggioni, 2006], and internal state values of long short-term memory [Kemeth et al., 2021].
@@ A.3 Hyperbolic Geometry
Hyperbolic geometry is a Riemannian manifold of a constant negative curvature [Lee, 2006]. There are four commonly used models for hyperbolic space: the Poincaré disk model, the Lorentz model, the Poincaré half-space model, and the Beltrami-Klein model. These models are equivalent and isometric, representing the same geometry, and distances can be translated without distortion [Ratcliffe et al., 1994]. In this work, we used the Poincaré half-space model to represent the continuous tree structure underlying the data because the exponential growth of the metric serves as a natural representation of the exponential propagation of the diffusion densities on a dyadic grid. Formally, the $n$-dimensional Poincaré half-space model $\mathbb{H}^{n}$ with constant negative curvature $-1$ is defined by $\mathbb{H}^{n}=\{\mathbf{x}\in\mathbb{R}^{n}\big{|}\mathbf{x}(n)>0\}$ with the Riemannian metric tensor $ds^{2}=(d\mathbf{x}^{2}(1)+d\mathbf{x}^{2}(2)+\ldots+d\mathbf{x}^{2}(n))/\mathbf{x}^{2}(n)$. The Riemannian distance for any $\bm{x},\bm{y}\in\mathbb{H}^{n}$ is defined by $d_{\mathbb{H}^{n}}(\mathbf{x},\mathbf{y})=2\sinh^{-1}(\left\lVert\mathbf{x}-\mathbf{y}\right\rVert_{2}/(2\sqrt{\mathbf{x}(n)\mathbf{y}(n)}))$. The distance function $d_{\mathbb{H}^{n}}(\cdot,\cdot)$ in the hyperbolic space $\mathbb{H}^{n}$ gives rise to two distinct types of geodesic paths [Ratcliffe et al., 1994]: straight lines that are parallel to the $n$-th axis, and semi-circles that are perpendicular to the boundary of the plane $\mathbb{R}^{n-1}$.
The Gromov product [Gromov, 1987] is used to define $\delta$-hyperbolic spaces in hyperbolic geometry.
@@
A metric space $(\mathcal{X},d_{\mathcal{X}})$ is $\delta$-hyperbolic space [Gromov, 1987] if there exists $\delta\geq 0$ such that for any four points $x_{1},x_{2},x_{3},x_{4}\in\mathcal{X}$, we have
@@
Definition A.2 (Gromov Product [Gromov, 1987]).
Let $(\mathcal{X},d_{\mathcal{X}})$ be a metric space. The Gromov product [Gromov, 1987] of any two points $x_{1},x_{2}\in\mathcal{X}$ w.r.t a reference point $x_{3}\in\mathcal{X}$ is defined by

### v3

@@ 5 Experimental Results
We apply the proposed method to a synthetic example and to word-document and single-cell RNA-sequencing (scRNA-seq) data. We use Word2vec (Mikolov et al., 2013) as word embedding vectors for the word-document data and Gene2vec (Du et al., 2019) as gene embedding vectors in scRNA-seq experiments. The implementation details are in App. D Additional experiments are in App. E.
@@ 5.1 Synthetic Word-Document Data Example
We compare the proposed TWD with existing TWDs, including (i) Quadtree (Indyk & Thaper, 2003) that constructs a random tree by a recursive division of hypercubes, (ii) Flowtree (Backurs et al., 2020) that computes the optimal flow of Quadtree, (iii) tree-sliced Wasserstein distance (TSWD) (Le et al., 2019) that computes an average of TWDs by random sampling trees, where the numbers of sampling are set to 1, 5, and 10, (iv) UltraTree (Chen et al., 2024) that constructs an ultrametric tree (Johnson, 1967) by minimizing OT regression cost, (v) weighted cluster TWD (WCTWD) (Yamada et al., 2022), (vi) weighted Quadtree TWD (WQTWD) (Yamada et al., 2022), and their sliced variants (vii) SWCTWD and (viii) SWQTWD (Yamada et al., 2022). We also compare to TWDs where the trees are constructed by (i) MST (Prim, 1957), (ii) Tree Representation (TR) (Sonthalia & Gilbert, 2020) through a divide-and-conquer approach, (iii) gradient-based hierarchical clustering (HC) in hyperbolic space (gHHC) (Monath et al., 2019), (iv) gradient-based Ultrametric Fitting (UltraFit) (Chierchia & Perret, 2019), and (v) HC by hyperbolic Dasgupta’s cost (HHC) (Chami et al., 2020). We refer to App. B for a detailed description of baselines and highlight the novel aspects of our TWD.
@@ 5.2 Real-World Word-Document Data Analysis
We demonstrate the power of the proposed TWD in real-world document classification tasks. The latent hierarchical structure of words (features) can be viewed as a node branching into related terms (Gollapalli & Caragea, 2014; Yurochkin et al., 2019; Chang & Blei, 2010), allowing for meaningful comparison between documents in OT distance (Kusner et al., 2015), as it measures how words and their associated meanings are transported from one document to another.
We test four word-document benchmarks in Kusner et al. (2015), including BBCSPORT, TWITTER, CLASSIC, and AMAZON. Detailed descriptions of these datasets are reported in App. D. We compare the proposed TWD to the same competing methods as in Sec. 5.1. In addition, we include the Word Mover’s Distance (WMD) (Kusner et al., 2015) that computes the OT distance between documents using Word2Vec (Mikolov et al., 2013). A dissimilarity classification based on our TWD and the baselines using the $k$NN classifier is applied. We use cross-validation with five trials, where the dataset is randomly divided into $70\%$ training set and $30\%$ testing set, following previous work in word-document data analysis (Kusner et al., 2015; Huang et al., 2016).
Tab. LABEL:tab:classification presents the document classification accuracy. The results of Quadtree, Flowtree, TSWD, and WMD are taken from Huang et al. (2016); Takezawa et al. (2021). Our method outperforms all the competing baselines on BBCSPORT, TWITTER, and AMAZON datasets, including the pre-trained WMD that is specifically designed for word-document data. On CLASSIC dataset, our proposed TWD yields the second-best classification accuracy and is competitive with WMD.
Table 1: Document and cell classification accuracy (the highest in bold and the second highest underlined). Results marked with are taken from Huang et al. (2016); Takezawa et al. (2021). The empirical comparisons to TWD methods demonstrate the advantages of our proposed approach, which effectively incorporates the latent feature hierarchy, leading to improved performance.
@@ 5.3 Single-Cell Gene Expression Data Analysis
We further demonstrate the advantages of our TWD in cell classification tasks on scRNA-seq data. The hierarchical structure of genes in scRNA-seq data (Tanay & Regev, 2017) are essential for analyzing complex cellular processes (Ding & Regev, 2021), especially for computing OT distance (Bellazzi et al., 2021; Tran et al., 2021; Bunne et al., 2023). For instance, in developmental biology, genes associated with stem cell differentiation follow a hierarchical pattern (Kang et al., 2012). Revealing the latent gene hierarchy is essential for accurately computing the TWD between cells, as it captures the underlying biological relationships and functional similarities.
We test two scRNA-seq datasets in Dumitrascu et al. (2021), including the Zeisel and CBMC. Detailed descriptions of the datasets are reported in App. D. We compare the proposed TWD to the same baselines as in Sec. 5.2. Instead of WMD, we include the Gene Mover’s Distance (GMD) (Bellazzi et al., 2021), a gene-based OT distance using Gene2Vec (Du et al., 2019). The cell classification tasks are performed in the same way as in Sec. 5.1 and Sec. 5.2. Tab. LABEL:tab:classification shows the mean and the standard deviation of the cell classification accuracy. Our TWD outperforms the competing methods by a large margin. This suggests that the proposed TWD effectively transforms the latent tree-like relationships in gene space, improving the cell classification accuracy compared to random sampling trees and trees constructed by graph-based and HC methods. Notably, our TWD differentiates the cell types better than pre-trained GMD, which uses gene proximity, while ours is geometrically intrinsic.
@@ 6 Conclusions
Ethics statement
This paper presents new machine learning techniques using publicly available and commonly used benchmarking datasets. These datasets are fully anonymized and do not contain any personal or sensitive information. We respect the privacy and integrity of all data, and we do not foresee any negative societal impact from our methods. We have no conflicts of interest to disclose.
Reproducibility Statement
We provide a detailed overview of our experimental setup, including the data splits, hyperparameter configurations, and the criteria for their selection. Additionally, we describe the type of classifier used and other relevant training details. For a complete explanation, please see App. D. Our code is available at https://github.com/ya-wei-eileen-lin/TWDwithLatentFeatureHierarchy.
@@ Acknowledgments
We thank the anonymous reviewers for their insightful feedback. We thank Gal Maman for proofreading an earlier version of the paper. The work of YEL and RT was supported by the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 802735-ERC-DIFFOP. The work of RC was supported by the US Air Force Office of Scientific Research (AFOSR MURI FA9550-21-1-0317). The work of GM was supported by NSF CCF-2217058.
@@ A.1 Wasserstein Distance
The application of the Wasserstein distance spans various fields, including enhancing generative models in machine learning (Arjovsky et al., 2017; Kolouri et al., 2018), aligning and comparing color distributions in computer graphics (Solomon et al., 2015; Lavenant et al., 2018), and serving as a robust similarity metric in data analysis (Bellazzi et al., 2021; Kusner et al., 2015; Yair et al., 2017).
Let $\mu_{1}$ and $\mu_{2}$ be two probability distributions on a measurable space $\Omega$ with a ground metric $d$, $\Pi(\mu_{1},\mu_{2})$ be the set of couplings $\pi$ on the product space $\Omega\times\Omega$ such that $\pi(B_{1}\times\Omega)=\mu_{1}(B_{1})$ and $\pi(\Omega\times B_{2})=\mu_{2}(B_{2})$ for any sets $B_{1},B_{2}\subset\Omega$. The Wasserstein distance (Villani, 2009) between $\mu_{1}$ and $\mu_{2}$ is defined by
@@
When replacing the $\inf$ in Eq. (4) with the $\arg\min$, the obtained solution $\pi$ is referred to as the Optimal Transport (OT) plan (Monge, 1781; Kantorovich, 1942). When computing the Wasserstein distance in Eq. (4) between two discrete probabilities with $m$ points, given by
@@
where $\mathbf{C}\in\mathbb{R}^{m\times m}$ is the ground pairwise distance matrix between the coordinates of the discrete probabilities, it can be computed in $O(m^{3}\log m)$ (Peyré et al., 2019), preventing its use of OT for large-scale datasets (Bonneel et al., 2011). Therefore, existing works have been proposed to reduce this complexity. Among them, a notable work is the Sinkhorn algorithm (Cuturi, 2013; Chizat et al., 2020), which approximates the OT in quadratic time.
@@ A.2 Diffusion Geometry
Diffusion geometry (Coifman & Lafon, 2006) is a mathematical framework designed to analyze high-dimensional data points $\{\mathbf{a}_{j}\in\mathbb{R}^{n}\}_{j=1}^{m}$ by effectively capturing their underlying geometric structures (Belkin & Niyogi, 2008). It focuses on analyzing the similarity between data points through the process of diffusion propagation. Specifically, the constructed diffusion operator $\mathbf{P}\in\mathbb{R}^{m\times m}$ derived from the observational data is associated with the heat kernel and the Laplacian of the underlying manifold (Belkin & Niyogi, 2008). Let $\mathbf{Q}(j,j^{\prime})=\exp\left(-d^{2}(j,j^{\prime})/\epsilon\right)$ be the $(j,j^{\prime})$th element of the pairwise affinity matrix between the data points. In the limit $m\rightarrow\infty$ and $\epsilon\rightarrow 0$, the diffusion operator $\mathbf{P}$ converges to the heat kernel $\exp(-\Delta)$, where $\Delta$ is the Laplace–Beltrami operator on the manifold (Coifman & Lafon, 2006). By applying eigendecomposition to the diffusion operator $\mathbf{P}$, we obtain a sequence of eigenpairs $\{\nu_{j},\bm{\varphi}_{j}\}_{j=1}^{m}$ with positive decreasing eigenvalues $01=\nu_{1}\geq\nu_{2}\geq\ldots\geq\nu_{m}>$ 0. The diffusion maps (DM) in $m^{\prime}<<m$ dimensions at the diffusion time $t$ embeds a point $\mathbf{a}_{j}$ into $\mathbb{R}^{m^{\prime}}$ by
@@
The diffusion operator has been shown to exhibit favorable convergence properties (Coifman & Lafon, 2006). In the limits, as the number of features $m$ increases toward infinity, the scaling parameter $\epsilon$ approaches zero, and the operator $\mathbf{P}^{t/\epsilon}$ converges pointwise to the Neumann heat kernel $\mathbf{H}_{t}=\exp(-t\Delta)$, associated with the underlying manifold, where $\Delta$ is the Laplace–Beltrami operator on the manifold. That is, the diffusion operator serves as a discrete approximation of the continuous heat kernel, thereby capturing the geometric structure of the manifold within a finite-dimensional framework (Coifman & Lafon, 2006; Singer, 2006; Belkin & Niyogi, 2008).
Note that the density-normalized affinity matrix was considered to mitigate the effect of non-uniform data sampling, which was proposed and shown theoretically in Coifman & Lafon (2006). In Section 3.1 in Coifman & Lafon (2006), the construction of a family of diffusion operators is introduced with a parameter $\alpha\in\mathbb{R}$ and the density $q$. In Section 3.4 in Coifman & Lafon (2006), it is shown that when $\alpha=1$ (in our case $\widehat{\mathbf{Q}}=\mathbf{D}^{-1}\mathbf{Q}\mathbf{D}^{-1}$), the diffusion operator approximates the heat kernel, and this approximation is independent of the density $q$. This technique has been further explored and utilized in (Mishne et al., 2019; Katz et al., 2020; Shnitzer et al., 2022; 2024; Lin et al., 2025).
Diffusion geometry has mainly been utilized in manifold learning, leading to the development of multi-scale, low-dimensional representations and informative distances (Nadler et al., 2005; Coifman & Maggioni, 2006). Its utility has been demonstrated across various applications spanning various fields, including nonlinear stochastic dynamical systems (Shnitzer et al., 2020), multi-view dimensionality reduction (Lindenbaum et al., 2020b), shape recognition (Bronstein et al., 2010), wavelet analysis (Coifman & Maggioni, 2006), internal state values of long short-term memory (Kemeth et al., 2021), and supervised and semi-supervised learning tools (Mendelman & Talmon, 2025).
@@ A.3 Hyperbolic Geometry
Hyperbolic geometry is a Riemannian manifold of a constant negative curvature (Lee, 2006). There are four commonly used models for hyperbolic space: the Poincaré disk model, the Lorentz model, the Poincaré half-space model, and the Beltrami-Klein model. These models are equivalent and isometric, representing the same geometry, and distances can be translated without distortion (Ratcliffe et al., 1994). In this work, we used the Poincaré half-space model to represent the continuous tree structure underlying the data because the exponential growth of the metric serves as a natural representation of the exponential propagation of the diffusion densities on a dyadic grid. Formally, the $n$-dimensional Poincaré half-space model $\mathbb{H}^{n}$ with constant negative curvature $-1$ is defined by $\mathbb{H}^{n}=\{\mathbf{x}\in\mathbb{R}^{n}\big{|}\mathbf{x}(n)>0\}$ with the Riemannian metric tensor $ds^{2}=(d\mathbf{x}^{2}(1)+d\mathbf{x}^{2}(2)+\ldots+d\mathbf{x}^{2}(n))/\mathbf{x}^{2}(n)$. The Riemannian distance for any $\bm{x},\bm{y}\in\mathbb{H}^{n}$ is defined by $d_{\mathbb{H}^{n}}(\mathbf{x},\mathbf{y})=2\sinh^{-1}(\left\lVert\mathbf{x}-\mathbf{y}\right\rVert_{2}/(2\sqrt{\mathbf{x}(n)\mathbf{y}(n)}))$. The distance function $d_{\mathbb{H}^{n}}(\cdot,\cdot)$ in the hyperbolic space $\mathbb{H}^{n}$ gives rise to two distinct types of geodesic paths (Ratcliffe et al., 1994): straight lines that are parallel to the $n$-th axis, and semi-circles that are perpendicular to the boundary of the plane $\mathbb{R}^{n-1}$.
The Gromov product (Gromov, 1987) is used to define $\delta$-hyperbolic spaces in hyperbolic geometry.
@@
A metric space $(\mathcal{X},d_{\mathcal{X}})$ is $\delta$-hyperbolic space (Gromov, 1987) if there exists $\delta\geq 0$ such that for any four points $x_{1},x_{2},x_{3},x_{4}\in\mathcal{X}$, we have
@@
Definition A.2 (Gromov Product (Gromov, 1987)).
Let $(\mathcal{X},d_{\mathcal{X}})$ be a metric space. The Gromov product (Gromov, 1987) of any two points $x_{1},x_{2}\in\mathcal{X}$ w.r.t a reference point $x_{3}\in\mathcal{X}$ is defined by